this is a d=rt problem Tina get a head start and leave work traveling at 45 mph three hour later Tim leaves work on the same road at 65 mph.how many mile will Tim pass Tina?
Answer by mananth(16946)

this is a d=rt problem Tina get a head start and leave work traveling at 45 mph three hour later Tim leaves work on the same road at 65 mph.how many mile will Tim pass Tina?
Tina 45 mph
leaves 3 hours early
so when Tim starts she is 45 * 3 = 135 miles ahead.
...
Tim speed = 65mph
..
the catch up speed = 65-45 20 mph
catch up distance = 135 miles
Time taken by Tim to catch up = 135/20 = 6.75 hours
..
Tim speed = 65 mph
time catch up = 6.75 hours
Distance he meets = 6.75 * 65=438.7 km
...
m.ananth@hotmail.ca